# Activation Addition: Steering Language Models Without Optimization

### v4

@@ Abstract
Reliably controlling the behavior of large language models is a pressing open problem. Existing methods include supervised finetuning, reinforcement learning from human feedback, prompt engineering and guided decoding. We instead investigate activation engineering: modifying activations at inference-time to predictably alter model behavior. We bias the forward pass with a ‘steering vector’ implicitly specified through natural language. Past work learned these steering vectors; our Activation Addition (ActAdd) method instead computes them by taking activation differences resulting from pairs of prompts. We demonstrate ActAdd on a range of LLMs (LLaMA-3, OPT, GPT-2, and GPT-J), obtaining SOTA on detoxification and negative-to-positive sentiment control. Our approach yields inference-time control over high-level properties of output like topic and sentiment while preserving performance on off-target tasks. ActAdd takes far less compute and implementation effort than finetuning or RLHF, allows users control through natural language, and its computational overhead (as a fraction of inference time) appears stable or improving over increasing model size.
@@ 1 Introduction
The success of large language models (LLMs) pretrained on massive corpora (Peters et al. 2018; Devlin et al. 2018; Radford et al. 2019; Brown et al. 2020) comes with a large caveat: our ability to control the output of these models remains partial & computationally expensive. Controls include supervised finetuning (Devlin et al. 2018), reinforcement learning from human feedback (RLHF) (Ziegler et al. 2019), prompt engineering (Radford et al. 2019), and guided decoding (Ghazvininejad et al. 2017; Gu et al. 2017).
We instead control LLMs by modifying their activations during the forward pass. We call this approach activation engineering, as pioneered by Subramani et al. 2022 and Hernandez et al. 2023. We go further by demonstrating natural-language control of activations and thereby output text. We use pairs of prompts to implicitly specify a direction in activation space which we scale to control the model (sentiment, truthfulness, but also topic and style; see Appendix Table 6). In contrast with past work (Shen et al. 2017; Dathathri et al. 2020; Subramani et al. 2022; Li et al. 2023b), ActAdd does not require labelled data (except prompts) nor backward passes to learn an encoder, controller, classifier or activation direction.
[…]
@@
We make four main contributions: 1) We find that combining forward passes (adding activation vectors) works well in LLMs despite the model not being trained for this; 2) We develop a lightweight control method, Activation Additions (ActAdd), which works at inference time and requires no optimization or labelled data. This is thus a promising direction for user control of LLMs; 3) We show that ActAdd preserves model performance, involves far less compute and implementation (vs finetuning and RLHF), scales naturally with model size, while being easier to implement; 4) We compare ActAdd  against several baselines and observe competitive results on sentiment control.
@@ 2 Related Work
Latent space arithmetic. Research in generative models for computer vision has long demonstrated the ability to steer image generation using derived vectors, including steering latent variables – most famously, intervening on a dimension that corresponds to smiles in images (Larsen et al. 2016; White 2016). Similarly, in the text domain, classic results on the word2vec embedding show that arithmetic on word vectors can capture some parts of semantic reasoning (for instance, analogies: Mikolov et al. 2013b; a). Our work differs in being computed (with forward passes) rather than learned (with backward passes); in operating on activation space, rather than embedding or weight space; and in presenting a natural-language user interface.
@@
  • Intervening on weights, as with supervised fine-tuning, RLHF, steerable layers, and weight editing (that is, targeted fine-tuning) (Ranzato et al. 2016; Ziegler et al. 2019; Dathathri et al. 2020; Meng et al. 2023; Ilharco et al. 2023). However, RLHF and weight editing have known side-effects on overall model performance (OpenAI 2023; Hase et al. 2023; Brown et al. 2023);
@@
  • Intervening on the prompt, as with automated prompt engineering (Shin et al. 2020; Zhou et al. 2022) and soft or continuous prompting (Lester et al. 2021; Li and Liang 2021;
  • Intervening on token embeddings, as with ‘soft prompting’ (Li and Liang 2021; Lester et al. 2021; Khashabi et al. 2022);
  • Intervening on activations, for instance by freezing the weights of the LLM and searching for a ‘steering vector’ of activations, e.g. using gradient descent (Subramani et al. 2022; Hernandez et al. 2023). These optimized extraction methods, which search for a steering vector, differ from extraction methods which directly compute it (present work and Li et al. 2023b). This paper elaborates on this latter family of methods.
Table 2: Locating our work in the steering vector literature.
@@
Steering vectors obtained via weights activations
@@
N/A ActAdd  (present work)
@@
Activation (or representation) engineering involves creating vectors of activations which cause desired changes to output text when added to the forward passes of a frozen LLM (Dathathri et al. 2020; Zou et al. 2023). Table 2 organizes prior work by intervention type.
An early antecedent is the Plug-and-Play Language Model of Dathathri et al. 2020. This uses a separate classifier (one classifier per attribute to steer towards) to perturb the model’s activations to generate text that accords more closely with the classifier’s target. Subramani et al. 2022 extract latent steering vectors from a frozen LLM, successfully discovering sentence-specific vectors which steer completions to near-perfect BLEU scores (i.e, control of the LLM’s generation) and unsupervised style transfer. However, the method requires running gradient descent for each new steering vector.
Independently, Li et al. 2023b developed a similar method which computes steering vectors (though not without some prior optimization). They use linear probe accuracy to find attention heads with different activation distributions for true and false statements (which requires labelled data, in this case TruthfulQA, to operationalize ‘truth’). The authors intervene on these heads to steer the model toward truthful output, where our experiments cover a range of goals without needing labelled data. In addition, ITI is repeated at each next-token prediction and requires dozens-to-hundreds of samples; suitable prompts can be found for ActAdd  in as few as 2 samples. Similar work, on ‘in-context vectors’ also followed ours (Liu et al. 2023).
@@
Hernandez et al. 2023 locate and edit an LLM’s knowledge through learning an encoding of facts in its activation space. Ablating attention heads can also be seen as activation engineering, though the technique is mostly used for model interpretation rather than steering per se (Michel et al. 2019; Olsson et al. 2022).
3 Methods
The Transformer.
We use decoder-only Transformer neural networks trained on a suitably large text corpus (Vaswani et al. 2017; Liu et al. 2018. The LLMs in this work are a stack of Transformer layers, each consisting of multi-head attention (MHA) and a feedforward network (FFN). We focus on its ‘residual streams’(Elhage et al. 2021), the sequences $(\mathbf{x}_{0},...,\mathbf{x}_{n})$ of token vectors processed by each layer. GPT-2-XL’s token vectors are 1600-dimensional (Radford et al. 2019). ActAdd manipulates the residual stream values $\mathbf{h}^{l}$ input to layer $l$. Following Elhage et al. 2021, we view each MHA block as adding an independent vector into that layer’s residual stream. Each layer performs MHA and FFN computations on $\mathbf{x}_{i}$, adding $\mathbf{x}_{i+1}$ to the stream. The final vector $\mathbf{x}_{n}$ in the stream can then be decoded into the next-token prediction. At inference-time, the residual stream is initialized $\mathbf{h}^{1}$ with the embedding of the tokenized prompt.
@@ 3 Methods
Activation addition. Our method takes a pair of natural-language prompts $(p_{+},p_{-})$, where $p_{+}$ represents the property we wish output text to emphasise and $p_{-}$ represents its opposite. $\mathbf{h}_{+}^{l}$ is the activation vector for the prompt $p_{+}$ at layer $l$. The difference $\mathbf{h}_{+}^{l}-\mathbf{h}_{-}^{l}$ is a new activation vector which (intuitively) captures the difference between a prompt with the property, and without it. This can be seen as an analogue of the ‘comparative preference statements’ used in recent formal logics (Kaci and Patel 2014). To obtain a steering vector, we perform a forward pass on each prompt, record the activations at the given location in each pass, take the difference $\mathbf{h}_{+}^{l}-\mathbf{h}_{-}^{l}$, and then finally rescale this difference in activations by an ‘injection coefficient’ $c$. To steer, we add the resulting activation vector to the input of layer $l$ and allow the forward pass to continue, and so obtain our steered output. (See Appendix C for implementation details.) $c$ represents intervention strength, since it multiplies the specified direction’s contribution to the residual stream and so the token processing of the rest of the forward pass. Absolute values between 3 and 15 are typical. Along with the target layer $l$, $c$ is a free parameter we select via grid search. (We find that, as expected from past work, intervening at middle layers is most effective (Subramani et al. 2022; Turner et al. 2023).)
Algorithm 1 and Figure 1 depict the resulting ActAdd  method. Appendix Figure 5 gives a figurative example of steering a model with ActAdd if that model had one-dimensional residual streams (rather than e.g. GPT-2-XL’s 1600 dimensions). The method differs from past work in being computed (with forward passes) rather than learned (with backward passes); in operating on activation space rather than embedding space or weight space; in presenting the user a natural-language interface; and in requiring as few as 2 samples in order to find an appropriate prompt pair (‘contrast pair’) to specify the steering direction. The contrast pair $(p^{+},p^{-})$ can be of arbitrary lengths. (One serious constraint is that the model must cache intermediate activations; see Bloom and Nanda 2022.) Further implementation details can be found in Appendix E. See also Appendix I for a partial variant of ActAdd. A runnable notebook can be found at tinyurl.com/actadd.
Interestingly, our steering vectors are not specified by taking the difference between desired outputs (e.g. “John married Jane" vs “John hired Jane"). Both prompts ($p_{+},p_{-}$) are (say) wedding-related: “I love talking about weddings" and “I hate talking about weddings". That this would lead to more steering toward the wedding topic is not a trivial fact, since both prompts talk about weddings and the instance of ‘wedding’ in $p_{-}$ becomes part of a vector which gets penalized. So ActAdd works above token-level. This is more analogous to ‘preference statements’ (Kaci and Patel 2014). Some desirable properties of a steering method include: effectiveness at steering output; generality (ability to affect any aspect of the output, e.g. sentiment, topic, style, knowledge, or role); ease of specification; low computational overhead; and no side effects on unsteered parts of the model. We design experiments to test ActAdd on these desiderata (see Results). Our initial experiments use GPT-2-XL (1.5B parameters, Radford et al. 2019), but the method generalizes and scales: see Figure 4, Table 13, and toxicity and sentiment experiments on OPT (Zhang et al. 2022b) and LLaMA-3 (Meta 2024).111 The full repo with all experiments can be found at https://zenodo.org/records/8215277. See Appendix E for reproducibility and F for replication.
3.1 Metrics
[…]
Shift in logprobs. To test if the intervention is affecting relevant tokens or reducing perplexity in some spurious way, we observe the shift in the distribution of token log probabilities. We do this by randomly sampling 500 documents from the above OpenWebText sample and recording the log-probabilities assigned by the baseline and ActAdded models. (This results in a dataset of  500k tokens, of which 29k are unique.) We then group by token, filter for tokens with $>\!\!20$ instances in the dataset, and calculate the mean $\mathcal{L}$ difference between the ActAdd  and baseline models. We display these as a Q-Q plot (Gnanadesikan and Wilk 1968, i.e. compare the distribution quantile values) and inspect outlier tokens.
Inference time premium. To measure how much overhead ActAdd adds to inference, (the percentage increase in time-to-complete one forward pass using ActAdd)  for different model sizes, we iterate over a list of models of different sizes and 10 random seeds. We obtain a baseline inference time for each (model, seed) pair through 100 repeated forward passes on a batch of random tokens (32 sequences of length 64). We obtain an ActAdd  inference time for each (model, seed) pair by running the previous method, augmented by a test ActAdd  contrast pair: ‘This is a test prompt.’ ($p_{+}$) and the empty string ($p_{-}$). Running a batch-of-2 forward pass on these gets us the activation addition tensor, which we add at layer 6. We take the mean inference time $\bar{t}$ over the 10 random seeds, and calculate the inference time premium as $\mathrm{premium}=\frac{\bar{t}_{\mathrm{ActAdd}}}{\bar{t}_{\mathrm{baseline}}}-1.$
Relevance steering. To demonstrate general topic steering, we test a range of subjects against a generic prompt, and use GPT-3.5 to score modified and unmodified model completions on whether they relate to the ActAdd topic (i.e. whether steering worked). The procedure: we generate 100 prompt completions from the unmodified model; iterate over a list of single-token topic ActAdds, applying each and generating 100 more completions; for each completion, we evaluate relevance to the topic using GPT-3.5. (The completions from the unmodified model are evaluated for relevance to each topic, to provide a relevance baseline; the modified-model completions are evaluated for relevance to the intended steering topic). Finally, we aggregate these evaluations for each (topic, coefficient) pair to get a mean relevance fraction, and subtract the baseline relevance.
Generation scoring To show the effect of different injection layers and give a sense of the reliability of ActAdd, we score the generations as follows. We generate a batch of completions for a specific prompt $p$, both with and without ActAdd, and computed average number of related words and fraction of completions with a related word over the resulting completions. The following settings are the only iteration run for this experiment: $p^{*}=\mathrm{`I\,went\,up\,to\,my\,friend\,and\,said}\textit{'},p^{+}=\mathrm{`weddings}\textit{'},$ ’ ,
$p_{-}=\mathrm{`\,\textit{'}},c=1.0,\,\mathrm{seed}=0.$ Completion length is 40 tokens with model sampling parameters: temperature $=1$, frequency penalty $=1$, top-P $=0.3$. For each setting, we compute statistics over 200 completions. Finally we sweep over all layers (i.e. for GPT-2, 1 – 48).
Preservation of performance, $P@K$. We also test that ActAdd  does not disrupt the model’s general knowledge (as some other steering methods do). We use ConceptNet from the LAMA benchmark, a general knowledge dataset (Petroni et al. 2019, $n=29774$ sentences, see Appendix Table 10). The test data involves prompting the model and filling the gap with the expected entity. The task is intended for both causal and masked models, so some examples are difficult for ‘causal’ models (like GPT-2) due to the extremely limited context.
Our evaluation procedure follows the original LAMA procedure: we load all sentences and extract the prompt and expected label. To simplify evaluation, we remove sentences with an expected label that tokenizes to more than one token. For each sentence, we run the model on its prompt with and without the wedding activation addition. $P@K$ is the probability that the expected label is among the model’s top-$K$ predicted tokens, conditioned on the prompt. We score the baseline and modified models by calculating mean $P@K$ values for a range of $K$. Finally we plot these for both modified and unmodified models over a range of $K$ values.
Sentiment steering. To evaluate sentiment we use the Stanford IMDb dataset (Maas et al., 2011), the goal of which is continuations with a sentiment opposite to that of the prompt. ‘Success’ is here the proportion of generated outputs with the desired sentiment, as classified by a model finetuned on sentiment data, SiEBERT (Hartmann et al. 2023b). For quality controls, we follow the conventional use of conditional perplexity to mark (dis)fluency, obtained using GPT-3 davinci-002 logprobs. We use cosine similarity between the prompt and continuation sentence embeddings to gauge the relevance of text in $[0,1]$. We evaluate sentiment changes from positive to negative and vice versa on a random subset $n=1000$ and repeat to obtain $p$-values. Code at tinyurl.com/actadd-sentiment.
Toxicity. We benchmark toxicity reduction by generating steered continuations from RealToxicityPrompts (Gehman et al., 2020). Following Pei et al. 2023 we use a random subset $n=1000$. We repeat this sampling 5 times to obtain $p$-values ($t$-test against SOTA), bolding rows which are better with $p<0.05$. For each continuation, we use the Perspective API to score toxicity (lower better). Code at tinyurl.com/actadd-toxicity.
4 Results

### v5

# Steering Language Models With Activation Engineering

@@ Abstract
Prompt engineering and finetuning aim to maximize language model performance on a given metric (like toxicity reduction). However, these methods do not fully elicit a model’s capabilities. To reduce this gap, we introduce activation engineering: the inference-time modification of activations in order to control (or steer) model outputs. Specifically, we introduce the Activation Addition (ActAdd) technique, which contrasts the intermediate activations on prompt pairs (such as “Love” versus “Hate”) to compute a steering vector (Subramani et al., 2022). By tactically adding in e.g. the “Love” -- “Hate” steering vector during the forward pass, we achieve SOTA on negative-to-positive sentiment shift and detoxification using models including LLaMA-3 and OPT. ActAdd yields inference-time control over high-level output properties (like topic and sentiment) while preserving performance on off-target tasks. ActAdd is lightweight: it does not require any machine optimization and works with a single pair of data points, which enables rapid iteration over steering. ActAdd demonstrates the power of activation engineering.
@@ 1 Introduction
LLMs contain hidden capabilities we do not know how to fully elicit (Korinek, 2023). Naively prompting a model with a question does not maximize the probability of the correct response. Carefully consider how prompting a model to think “step-by-step” (Wei et al., 2022) massively improves performance on a range of benchmarks. Similarly, “few-shot” prompting a model with correct answers to unrelated in-distribution questions allows “in-context learning” for e.g. stronger performance on NLP tasks (Brown et al., 2020). Importantly, these interventions do not supply the LLM with extra task-relevant information or update the algorithm implemented by the LLM’s computational graph. Even though the model is initially able to score higher on these benchmarks, those capabilities do not emerge without a specific intervention. We therefore hypothesize the presence of an elicitation overhang: we do not know how to elicit all relevant abilities and information from frontier models.
Prompt engineering is the most obvious way to steer a model, but prompting has limited reliability (Ye & Durrett, 2022; Wang et al., 2024). Therefore, to reduce the elicitation overhang, we explore a new modality for steering language model outputs. By strategically perturbing activations during the forward pass, we hope to more reliably and effectively steer models compared to prompt engineering. We call this methodology activation engineering.
We suspect that compared to prompt engineering, activation engineering can elicit a wider range of model capabilities. Consider, for example, a model optimized to imitate the text outputs of eloquent poets and awkward mathematicians. The model may contain the internal mechanisms required to output text which is both eloquent and mathematical. However, if the model is an accurate estimator of the training distribution, it will (correctly) assign low probability to eloquent mathematical prose. Because nothing in the training data was both eloquent and mathematical, there may exist no prompt which elicits mathematical prose. In contrast, activation engineering might be able to simultaneously activate the circuitry for eloquent speech and for mathematical content.
To demonstrate the power of activation engineering, we introduce Activation Addition (ActAdd). Suppose we want to achieve negative-to-positive sentiment control (Li et al., 2018; Dathathri et al., 2020). To achieve this, ActAdd first compares the model’s activations on a contrast pair of prompts, such as the prompts “Love” and “Hate.” The otherwise-similar prompts differ along the target dimension of sentiment. ActAdd then computes the difference of these activations in order to compute steering vectors. These vectors act like “virtual bias terms” because ActAdd directly adds the steering vectors to the forward pass at inference time. By shifting the inference-time activations along the direction of the steering vector, ActAdd steers the model to generate positive sentiment completions (Table 1).
Table 1: The impact of ActAdd. The steering vectors are computed from (“Love” - “Hate”) and (“I talk about weddings constantly” - “I do not talk about weddings constantly”). Appendix Table 6 shows more examples.
@@
Contributions.
We unify past literature on related topics to introduce activation engineering. To better elicit the full capabilities of models, we introduce the ActAdd steering method, which achieves SOTA on toxicity reduction and sentiment control. We thoroughly test the steered models to verify the preservation of their general capabilities. We therefore show the promise of ActAdd as an effective and cheap method for steering LLM outputs.
@@ 2 Related Work
Latent space arithmetic. Research in generative models for computer vision has long demonstrated the ability to steer image generation using derived vectors, including steering latent variables – most famously, shifting activations along a direction that corresponds to smiles in images (Larsen et al. 2016; White 2016). Similarly, in the text domain, classic results on the word2vec embedding show that arithmetic on word vectors can capture some parts of semantic reasoning (for instance, analogies: Mikolov et al. 2013b; a). Our work focuses on steering generative language models.
@@
  • Intervening on weights, as with supervised finetuning, RLHF, steerable layers, and weight editing (that is, targeted fine-tuning) (Ranzato et al. 2016; Ziegler et al. 2019; Dathathri et al. 2020; Meng et al. 2023; Ilharco et al. 2023). However, naive RLHF, finetuning, and weight editing have known side-effects on overall model performance (Hase et al. 2023; Qi et al. 2023; Brown et al. 2023);
@@
  • Intervening on the prompt, as with automated prompt engineering (Shin et al. 2020; Zhou et al. 2022);
  • Intervening on token embeddings, as with ‘soft prompting’ (Li & Liang 2021; Lester et al. 2021; Khashabi et al. 2022);
  • Intervening on activations, for instance by freezing the weights of the LLM and searching for a ‘steering vector’ of activations, e.g. using gradient descent (Subramani et al. 2022; Hernandez et al. 2023). These optimized extraction methods, which search for a steering vector, differ from extraction methods which directly compute it (present work and Li et al. 2023b). In our work, we do not use gradient descent or other optimization methods.
Table 2: Locating our work in the steering literature.
@@
Intervention vectors obtained via weights activations
@@
N/A
ActAdd (present work),
Li et al.,2023b
@@
Activation engineering involves creating vectors of activations which cause desired changes to output text when added to the forward passes of a frozen LLM (Dathathri et al. 2020). Table 2 organizes prior work by intervention type.
An early antecedent is the Plug-and-Play Language Model of Dathathri et al. 2020. This uses a separate classifier (one classifier per attribute to steer towards) to perturb the model’s activations to generate text that accords more closely with the classifier’s target. Subramani et al. 2022 extract latent steering vectors from a frozen LLM, successfully discovering sentence-specific vectors which steer completions to near-perfect BLEU scores (i.e, control of the LLM’s generation) and unsupervised style transfer. However, the method requires running gradient descent for each new steering vector. Hernandez et al. 2023 locate and edit an LLM’s knowledge through learning an encoding of facts in its activation space. Ablating attention heads can also be seen as activation engineering, though the technique is mostly used for model interpretation rather than steering (Michel et al. 2019; Olsson et al. 2022).
Independently, Li et al. 2023b developed a similar method called ITI which computes steering vectors which are selectively applied according to trained linear probes. They use these probes to find attention heads with different activation distributions for true and false statements. They steer the model toward truthful outputs, where our experiments cover a range of goals. In addition, ITI adds the same steering vector at all sequence positions during inference and ITI requires dozens of samples. In contrast, ActAdd we add steering vectors to a subset of sequence positions and require as few as 2 samples. Similar work on ‘in-context vectors’ also followed ours (Liu et al. 2023). Lastly, Zou et al. 2023’s “representation engineering” also followed our work. They develop a range of techniques for deriving steering vectors and for steering models using activation-space edits and optimization. In comparison to Zou et al. 2023, we steer different models (LLaMA-3, OPT, GPT-2, and GPT-J) on different tasks (detoxification and sentiment control).
3 How Activation Addition works
We use decoder-only Transformer neural networks (Vaswani et al. 2017). The LLMs in this work contain a stack of Transformer layers, each consisting of multi-head attention (MHA) and a feedforward network (FFN). We focus on its “residual streams” (Elhage et al. 2021), the sequences $(\mathbf{x}_{0},...,\mathbf{x}_{n})$ of intermediate activation vectors processed by each layer. ActAdd manipulates the residual stream values $\mathbf{h}^{l}$ input to layer $l$. Each layer performs MHA and FFN computations on $\mathbf{x}_{i}$, adding $\mathbf{x}_{i+1}$ to the stream. The final vector $\mathbf{x}_{n}$ in the stream can then be decoded into the next-token prediction. At inference time, the residual stream is initialized $\mathbf{h}^{1}$ with the embedding of the tokenized prompt.
Activation addition. Our method takes a pair of natural-language prompts $(p_{+},p_{-})$, where $p_{+}$ represents the property we wish output text to emphasise (e.g. love) and $p_{-}$ represents its opposite (e.g. hate or indifference). $\mathbf{h}_{+}^{l}$ is the activation vector for the prompt $p_{+}$ at layer $l$. The difference $\mathbf{h}_{+}^{l}-\mathbf{h}_{-}^{l}$ is a new activation vector which (intuitively) captures the difference between a prompt with the target property, and a prompt without it. The steering vector is computed before inference time.
@@ 3 How Activation Addition works
To obtain a steering vector, we perform a forward pass on each prompt, record the activations at the given location in each pass, take the difference $\mathbf{h}_{+}^{l}-\mathbf{h}_{-}^{l}$, and then finally rescale this difference in activations by an ‘injection coefficient’ $c$. To steer, we add the resulting activation vector to the input of layer $l$ and allow the forward pass to continue, and so obtain our steered output. [^1] See Appendix C for implementation details. $c$ represents the intervention strength, since it multiplies the steering vector’s contribution to the residual stream. [^2] It’s typical for the intervention strength $c$ to have a magnitude less than 15. We perform hyperparameter tuning to select $c$ and also the injection layer $l$. As expected from past work (Subramani et al. 2022; Mini et al. 2023), intervening at the middle layers is most effective.
Algorithm 1 and Figure 1 depict the resulting ActAdd method. In the appendix, Figure 6 illustrates a figurative example of steering a model with ActAdd if that model had one-dimensional residual streams (rather than e.g. GPT-2-XL’s 1600 dimensions). A runnable notebook can be found at tinyurl.com/actadd.
We test whether 1) steering vectors are effective at eliciting the desired behavioral shift, and 2) whether they preserve the general capabilities of the model. We run perplexity-based experiments on GPT-2-XL (1.5B parameters, Radford et al. 2019). We then run toxicity and sentiment experiments on OPT (Zhang et al. 2022b) and LLaMA-3 (Meta 2024).
4 Results: Activation Addition works